Two cars are moving at a constant velocity when they are involved in a collision. What kind of forces are acting to each other during the collision? After the collision both separate from each other. One care comes to rest after 0.50s with an average acceleration of 80.0 m/s^2 in the direction opposite that of the car's. What is its speed the speed, in km/h, of the car right after the collision? what rule can be used if you want to determine their speeds right before and after the collision of two cars?
Answer #1

During the collsion, the fricitonal force between the cars exist. In this case, there is no dissipative force acting during the collision.

From the kinematic equation of motion

v = u + at = 0 + ( 80 ) ( 0.50 s ) = 40 m/s = 40 ( 3600 / 1000 ) km/h = 144 km /h

-----------------------------------------------------------------------------------------------------------------------------------

In the case of elastic collison, the law of conservation of momentum and the law of conservation of kinetic energy holds good. If the collision is inelastic, only momnetum of the system is conserved, the kinetic energy not conserved.